If the probability of winning a match is 0.7 then what is the value of variance?
Show answerHide answer
Correct answer: (C) 0.21
Explanation
For a single trial (Bernoulli), variance = p × q = 0.7 × 0.3 = 0.21.
Correct answer: (C) 0.21
Explanation
For a single trial (Bernoulli), variance = p × q = 0.7 × 0.3 = 0.21.
This is question 86 of DSSSB TGT Computer Science PYQ (Technical), 7 Aug 2021, Shift 2. See all questions of this paper with answers, or attempt the paper as a timed online test to get your score and all-India rank.